OpenStudy (anonymous):

How many 3-digit passwords can be formed from the digits 0 through 9?

OpenStudy (anonymous):

answers 729 900 1,000

OpenStudy (mathteacher1729):

This is asking "how many numbers are there from 0 to 999 ? " :)

OpenStudy (mathteacher1729):

(remember to include 0)

OpenStudy (anonymous):

\[9 \times 10\times 10 =900 \]

OpenStudy (anonymous):

why not 1,000?

Parth (parthkohli):

9 slots for first digit. 10 for second. 10 for third. 9 * 10 * 10

OpenStudy (anonymous):

like what tyler said?

OpenStudy (anonymous):

so what one

Parth (parthkohli):

What is 9 * 10 * 10?

Parth (parthkohli):

The first digit cannot be 0, because then that'll become a 2 digit number

OpenStudy (anonymous):

so 900

Parth (parthkohli):

Yes.

OpenStudy (mathteacher1729):

@FoolForMath that is incorrect. They want all the passwords from 000 through 999 So that's 10 choices for each of the three places. 10^3 = 1000. :) or just list them all 000 001 002 003 ... 997 998 999 Which is 1,000 total codes.

OpenStudy (anonymous):

Thanks math teacher!

OpenStudy (anonymous):

000 is not 3 digit

OpenStudy (anonymous):

Hm, abstruse, depends on the implementation.

OpenStudy (anonymous):

It is not clarified whether 000 is 3 digit or 1 digit.

Parth (parthkohli):

I'm not sure what 001....099 are

Parth (parthkohli):

lo

OpenStudy (mathteacher1729):

There are two ways to interpret this problem. 1) "Three digit codes" means "numbers from 100 to 999". 2) "Three digit codes" means "numbers from 000 to 999".

OpenStudy (anonymous):

Yes.

Parth (parthkohli):

If either 1000 or 900 weren't a choice, then it'd have been solved easily.

Parth (parthkohli):

They have made it unclear and confusing.

OpenStudy (anonymous):

Poorly worded problem.

OpenStudy (mathteacher1729):

I chose to interpret it as option 2 because it specifically said codes, and one can set a combination lock or 3-digit passcode as a string of zeros (or 001 etc) if one likes. :)

Parth (parthkohli):

Nobody can be sure. You may go for 1000 or may not.

OpenStudy (anonymous):

ill choose 1000

OpenStudy (anonymous):

Yes, I agree with @mathteacher1729, but it depends solely on the implementation.

Parth (parthkohli):

I go with Mr. Moderator.

OpenStudy (mathteacher1729):

This reminds me of a scene from Spaceballs... http://youtu.be/a6iW-8xPw3k

OpenStudy (anonymous):

And in such cases, I will choose not answer this question.